Focus on Concepts, Question 15 Blood flows through a section of a horizontal artery that is partially blocked by a deposit alBlood flows through a section of a horizontal artery that is partially blocked by a deposit along the artery wall. What is the speed of the blood in the partially blocked region?
Answer #1

Volume flow rate at both section is same,

A1*v1 = A2*v2 where A1,A2 are area area of cross sections and v1,v2 are respective velocity

pi*2^2*0.40 = pi*1^2*v2

v2 =0.4*4 = 1.6 m/s^2

Fourth option is correct.